Find the area of the triangle formed by the line 2x – y + 10 =0 and the axes.
Answer by Alan3354(69443)

Find the area of the triangle formed by the line 2x – y + 10 =0 and the axes.
------
Find the x and y intercepts.
2x – y + 10 =0
2x – y = -10
x = 0, y = 10 --> (0,10)
y = 0, x = -5 --> (-5,0)
-----
Area = b*h/2 = 10*5/2
Area = 25 sq units